OpenStudy (anonymous):

Let u = <-6, 3>, v = <1, 9>. Find u - v.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

u - v is the same as u + (-v)

jimthompson5910 (jim_thompson5910):

u + (-v) is the same as u + (-1)*v

OpenStudy (anonymous):

how would i start

OpenStudy (yanasidlinskiy):

Just add the coordinates together.

OpenStudy (anonymous):

but thats not one of the answers

OpenStudy (anonymous):

<-15, 2> <-9,-8> <-5, 12> <-7, -6>

jimthompson5910 (jim_thompson5910):

you need to compute (-1)*v first

jimthompson5910 (jim_thompson5910):

then you can add

OpenStudy (anonymous):

i so far have -6,3 and -1,-9

OpenStudy (yanasidlinskiy):

I thought you would just plug it in as they give you the eqation: For the first you plug in -6-3 = -9 And then for the second one do the same thing....1-9 = -8 Your coordinates would be -9,-8 Isn't that how you are suppose to do it?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Now add those vectors melacho

OpenStudy (anonymous):

-6 +3?

jimthompson5910 (jim_thompson5910):

no you add the corresponding components so -6 + (-1) = ??? 3 + (-9) = ???

OpenStudy (anonymous):

<-7, -6>

jimthompson5910 (jim_thompson5910):

correct another way is you subtract the corresponding coordinates: u - v = <-6, 3> - <1, -9> u - v = <-6 - 1, 3 - 9> u - v = <-7, -6>

jimthompson5910 (jim_thompson5910):

either way, you get the same answer

OpenStudy (anonymous):

THANK YOU SO MUCH♥

jimthompson5910 (jim_thompson5910):

np